The Gauteng chamber of business conducted a survey amongst 17 furniture retailers to identify the percentage of bad debts in each company’s debtors’ book. The bad debts percentages are as follows:

 

2.2, 4.7, 6.3, 5.8, 5.7, 7.2, 2.6, 2.4, 6.1, 6.8, 2.2, 5.7, 3.4, 6.6, 1.8, 4.4, 5.4


 

Calculate the Pearson Coefficient of skewness coefficient for percentage of bad debts. Is the data skewed?



1
Expert's answer
2021-05-28T09:49:10-0400

Pearson Coefficient of skewness coefficient


For a given set of values

2.2, 4.7, 6.3, 5.8, 5.7, 7.2, 2.6, 2.4, 6.1, 6.8, 2.2, 5.7, 3.4, 6.6, 1.8, 4.4, 5.4


Mean = (2.2 + 4.7 + 6.3 + 5.8 + 5.7 + 7.2 + 2.6 + 2.4 + 6.1 + 6.8 + 2.2 + 5.7 + 3.4 + 6.6 + 1.8 + 4.4 + 5.4)/17

=79.3/17

=4.66470588235



Median =5.4


Variance (σ2)

"=(2.2 - 4.66)^2+ (4.7 - 4.66)^2+ (6.3 - 4.66)^2 + (5.8- 4.66)^2+\\(5.7- 4.66)^2+(7.2- 4.66)^2+ (2.6 - 4.66)^2+ (2.4 - 4.66)^2 +\\ ( 6.1- 4.66)^2+ (6.8 - 4.66)^2+ ( 2.2 - 4.66)^2 
+(5.7- 4.66)^2+\\ ( 3.4 - 4.66)^2+ (6.6 - 4.66)^2 + ( 1.8- 4.66)^2+ (4.4 - 4.66)^2+ ( 5.4 - 4.66)^2] / 17


\\
=1.79638237"

standard deviation (σ)


"=\sqrt{3.22698961925}\\
=1.79638237"


Pearson's Coefficient of Skewness coefficient



"=\frac{3 (mean- median)}{σ}\\
=\frac{3 (4.66470588235- 5.4)}{1.79638237}\\
=-1.228"


If skewness is negative, the data are negatively skewed or skewed left.